OpenStudy (anonymous):

The length of the hypotenuse of a right triangle is 16 inches. If the length of one leg is 5 inches, what is the approximate length of the other leg? 10.5 inches 11.0 inches 15.2 inches 16.8 inches

OpenStudy (dayakar):

second leg square = hypotenuse square- first leg square = 16^2-5^2 =256-25 =231

OpenStudy (dayakar):

\[second leg = \sqrt{231}\]

OpenStudy (dayakar):

\[\approx=15.2 inches\]

OpenStudy (dayakar):

\[\approx15.2\]

OpenStudy (dayakar):

is it clear